How would you solve the equation .25^x-0.5=2 ?
Answer by helpnalgebra(91)

25^(x)-0.5=2
25^(x)=0.5+2
25^(x)=2.5
ln(25^(x))=ln(2.5)
xln(25)=ln(2.5)
xln(25)=(0.92)
(xln(25))/(ln(25))=(0.92)/(ln(25))
x=(0.92)/(ln(25))_xAPPR0.28 answer